OpenStudy (ashy98):

Which of the following is not a way to represent the solution of the inequality 5x + 10 is greater than or equal to 7x + 2x - 2? A.) A number line with a closed circle on 3 and shading to the right B.) A number line with a closed circle on 3 and shading to the left C.) 3 is greater than or equal x D.) x is greater than or equal 3

OpenStudy (ashy98):

@Haseeb96

OpenStudy (haseeb96):

what do u think which one is correct?

OpenStudy (ashy98):

my bad its A

OpenStudy (anonymous):

5x+10 >=7x+2x-2 5x+10>=9x-2 adding +2 for both sides 5x+10+2>=9x-2+2 5x+12>=9x adding -5x for both sides 5x+12-5x>=9x-5x 12>=4x divide both sides by 4 3>=x so 3 is greater than or equal x

OpenStudy (anonymous):

b and c are both correct just different ways of saying the same thing BUT you are asked for NOT a way so A and D is the answer

OpenStudy (ashy98):

the answer was a